OpenStudy (anonymous):

Medal and fan for help please help The sets of numbers 7, 24, 25 and 9, 40, 41 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples! (written response)

OpenStudy (anonymous):

@sleepyhead314 i need help

OpenStudy (anonymous):

@CausticSyndicalist please help when you can

OpenStudy (sleepyhead314):

ummm try to use the Pythagorean theorem? a^2 + b^2 = c^2

OpenStudy (anonymous):

ok so it would be 7^2+ 24^2= 25^2 and 9^2+ 40^2= 41^2

OpenStudy (sleepyhead314):

yes and 7^2+ 24^2= 25^2 simplifies to 625 = 625 which is a true statement

OpenStudy (anonymous):

and then it would become 49+576=625 81+1600=1681

OpenStudy (anonymous):

ok so 625=615 and 1681=1681

OpenStudy (sleepyhead314):

yep... but I'm not exactly sure what the question is asking :/ http://en.wikipedia.org/wiki/Pythagorean_triple

OpenStudy (anonymous):

ok thanks for your help old friend i will message you if i need more help but i think i got it from here thanks.

OpenStudy (sleepyhead314):

glad I could help :)